# Any-Precision LLM: Low-Cost Deployment of Multiple, Different-Sized LLMs

### v3

@@ 2.2 LLM Quantization
This section discusses recent advancements in LLM quantization with a specific emphasis on weight-only, post-training quantization (PTQ). In LLMs, there is a pronounced shift towards quantizing only the weights, as the dominant bottleneck in inference throughput is the memory constraint imposed by the size of weight parameters, rather than computational requirements. Moreover, Post-Training Quantization (PTQ) has become a favored method for quantizing LLMs due to its practicality. Although Quantization-Aware Training (QAT) typically yields superior performance, its high training expense frequently renders it impractical (Liu et al., 2023; Dettmers et al., 2023a; Kim et al., 2023a).
@@
While the aforementioned methods employ uniform quantization, non-uniform quantization may be a more effective alternative as it better captures the weight distributions (Gholami et al., 2021). SqueezeLLM (Kim et al., 2023b) proposes a clustering-based LLM quantization that considers the sensitivity of each weight. Somewhat orthogonal to the aforementioned studies that primarily focus on rounding schemes, there are also proposals to use mixed precision as a way of allocating more bits to sensitive weights (Kim et al., 2023b; Dettmers et al., 2023b; Lee et al., 2023).
@@ 3.1 Need for Deploying Multiple, Different-sized LLMs
Deploying a set of different-sized LLMs provides significant practical advantages. It enhances user experience by effectively handling queries with varying latency requirements. Depending on the specific needs of users and applications, some queries require quick responses, while others can tolerate slower response times. Latency requirements become even more diverse when serving multiple different tasks concurrently, a common use case of LLMs. For instance, queries for interactive tasks like chatbots are mostly latency-sensitive, while tasks like document analysis, often handled in the background, allow for more relaxed response times.
@@ 4.2 Non-uniform Quantization-based Incremental Upscaling
In contrast to clustering-based methods, state-of-the-art uniform quantization methods often involve complicated mechanisms such as weight reconstruction (Frantar et al., 2023) and per-channel scaling (Lin et al., 2023) that are not readily compatible with incremental upscaling. For example, Table 2 presents the results of applying incremental upscaling to two state-of-the-art uniform methods (AWQ, GPTQ), using a 3-bit model as the seed model. The upscaled models consistently underperform the independently quantized models, showing unacceptable degradation at low bit-widths such as 4. We provide an in-depth discussion on the challenges of applying incremental upscaling on existing uniform quantization methods, along with additional experimental results, in Appendix F.
@@ 5.3 GPU Kernel Optimization
Efficient Bit-transpose. Dequantizing weights represented as bitplanes requires a large number of bitwise operations, becoming a potential bottleneck. The bit-transpose step (Step 2 in Figure 6), in particular, is the main overhead. Even an optimized bit-transpose algorithm requires 38 bitwise operations just to process an 8$\times$8 bit matrix (Warren, 2012).
@@ 6.1 Any-Precision Quantization Results
Methodology. We evaluate 4 to 8-bit models obtained through incremental upscaling, using a 3-bit SqueezeLLM model as the seed model. The results are compared with the 4 to 8-bit SqueezeLLM models. We benchmark our method on LLaMA-2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and three OPT models (6.7B, 2.7B, 1.3B) (Zhang et al., 2022). We evaluate the models with two metrics: perplexity on three datasets (WikiText2 (Merity et al., 2016), PTB (Marcus et al., 1994), C4 (Raffel et al., 2023)) and zero-shot accuracies on five tasks (ARC-easy/challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Tata & Patel, 2003), WinoGrande (Sakaguchi et al., 2021)). Experimental details are in Appendix A.
@@
Results. Table 3 shows the results. The upscaled models (SqLLM+IU) nearly match the preplexity of independently quantized models (SqLLM) across various models, datasets, and bit-widths. Except in two cases (OPT-6.7B, OPT-1.3B 4-bit on PTB), the perplexity increase is negligible (< 0.1). This demonstrates that with just the single $n$-bit parent model generated through the incremental upscaling process, we can utilize the full range of 3 to $n$-bit models, all achieving the state-of-the-art quality at their respective bit-widths. Zero-shot task results are in Appendix B.
Runtime of Incremental Upscaling. Although our incremental upscaling process consists of a large number of K-means clustering tasks, all tasks are trivially parallelizable. On an Intel i9-13900K CPU with 24 cores, it takes only about 2 minutes to produce 4, 5, …, 8-bit Llama-2-7B models from a 3-bit seed model. More results on the runtime of the incremental upscaling process are in Appendix C.
@@ 6.2 Kernel Microbenchmarks
Methodology. We benchmark the matrix-vector multiplication performance of our kernel on the three matrix dimensions used in Llama-2-7B, and compare with the existing kernel for non-uniform quantization proposed in SqueezeLLM. Note that the SqueezeLLM kernel only supports 3 and 4 bits. We conduct experiments on three GPUs of varying scales: RTX 4090 (desktop), RTX 4070 Laptop (laptop), and Jetson AGX Orin 64 GB (mobile). We use NVIDIA Nsight Compute to measure latency.
Results. Table 4 shows the results. Our kernel consistently achieves low latency for various matrix sizes and platforms, showing a near-linear improvement against the cuBLAS FP16 baseline as the bit-width decreases. Compared to the SqueezeLLM kernel, our kernel performs on par on the RTX cards, and goes on to exhibit substantial improvements on Jetson. Note that our kernel achieves the performance despite adopting the bitplane-based weight representation to support any-precision, which entails non-trivial engineering challenges such as addressing the higher dequantization overhead. Additional results with a wider range of matrix dimensions, comparison against uniform quantization kernels, and matrix-matrix multiplication performance for non-single batch scenarios are in Appendix D.
@@
6.3 End-to-end Performance
We evaluate the end-to-end inference performance of our engine by integrating it with TensorRT-LLM (NVIDIA, ). Figure  11 shows the throughput of Llama-2-7B in generating 128 and 1024 tokens on RTX 4090 and RTX 4070 Laptop. TensorRT-LLM currently lacks support for Jetson. The enhanced kernel performance effectively translates into an end-to-end speedup. A slightly lower speedup with a longer sequence length (1024) is due to the increased overhead of attention, which is not accelerated by weight quantization. Results for the other models are in Appendix E.
@@ 7 Conclusion
We make a case for any-precision LLM, that enables memory-efficient and cost-effective deployment of multiple, different-sized LLMs. We suggest a lightweight method for any-precision quantization of LLM, along with a specialized software engine to fully leverage its benefits.
@@ A.3 Zero-shot Tasks
Appendix B Results on Zero-shot Tasks
The data in Table 7 presents the zero-shot task accuracies of the standard SqueezeLLM (SqLLM) with its incremental upscaling counterpart (SqLLM+IU) starting from a 3-bit seed model. The comparison includes the average accuracies across tasks for each setup. The results indicate that the SqLLM models with incremental upscaling (SqLLM+IU) achieve accuracies very similar to those of the independently quantized SqLLM models across a range of tasks, models, and bit-widths.
Appendix C Runtime of Incremental Upscaling
Table 5: Runtime of the incremental upscaling process on SqueezeLLM. The process begins with a 3-bit seed model and progresses up to the final, 8-bit parent model. Measured on an Intel Core i9-13900K machine.
Llama-2-7B Mistral-7B OPT-6.7B OPT-2.7B OPT-1.3B
125 s 140 s 133 s 54 s 27 s
Table 5 shows the runtime of the incremental upscaling process on SqueezeLLM for five different models.
Appendix D Additional Kernel Microbenchmark Results
D.1 Kernel Latency on Various Matrix Sizes
@@
D.2 Comparison with Kernels for Uniform Quantization
@@
D.3 Matrix-Matrix Multiplication Performance
While single batch would be the predominant choice for our target setting of a single user running a model on personal devices, there are also some use cases where a small batch size, such as 2 or 4, is employed. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023d) and parallelization of multiple function calls for complex reasoning tasks (Kim et al., 2023e) are examples.
@@ D.3 Matrix-Matrix Multiplication Performance
Appendix E Additional End-to-End Performance Evaluation Results
@@
Appendix F Incremental Upscaling with Uniform Quantization Methods
@@
In this section, we elaborate on our endeavor to apply incremental upscaling to the two state-of-the-art uniform quantization methods (GPTQ, AWQ), which ultimately results in failure. Figure 13 illustrates the process of incremental upscaling with uniform quantization methods. Upscaling on uniform quantization methods requires that each bin in the quantization grid be divided equally; this is in contrast with non-uniform quantization, where each bin in the quantization grid can be divided into two new bins of arbitrary sizes. Naturally, each bin’s midpoint becomes its representative value. With these premises, Appendix F.1 and  F.2 discuss issues with upscaling GPTQ and AWQ, respectively.
F.1 Incremental Upscaling with GPTQ
@@
F.2 Incremental Upscaling with AWQ

### v4

@@ 2.2 LLM Quantization
This section discusses recent advancements in LLM quantization with a specific emphasis on weight-only, post-training quantization (PTQ). In LLMs, there is a pronounced shift towards quantizing only the weights, as the dominant bottleneck in inference throughput is the memory constraint imposed by the size of weight parameters, rather than computational requirements. Moreover, Post-Training Quantization (PTQ) has become a favored method for quantizing LLMs due to its practicality. Although Quantization-Aware Training (QAT) typically yields superior performance, its high training expense frequently renders it impractical (Liu et al., 2023; Dettmers et al., 2023; Kim et al., 2023a).
@@
While the aforementioned methods employ uniform quantization, non-uniform quantization may be a more effective alternative as it better captures the weight distributions (Gholami et al., 2021). SqueezeLLM (Kim et al., 2023b) proposes a clustering-based LLM quantization that considers the sensitivity of each weight. Somewhat orthogonal to the aforementioned studies that primarily focus on rounding schemes, there are also proposals to use mixed precision as a way of allocating more bits to sensitive weights (Kim et al., 2023b; Dettmers et al., 2024; Lee et al., 2024).
@@ 3.1 Need for Deploying Multiple, Different-sized LLMs
Deploying a set of different-sized LLMs provides significant practical advantages. It enhances user experience by effectively handling queries with varying latency requirements. Depending on the specific needs of users and applications, some queries require quick responses, while others can tolerate slower response times. Latency requirements become even more diverse when serving multiple different tasks concurrently, a common use case of LLMs. For instance, queries for interactive tasks like chatbots are mostly latency-sensitive, while tasks like document analysis, often handled in the background, allow for more relaxed response times. In fact, the approach of adaptively selecting different DNN models with various accuracy-latency trade-offs has been widely studied as a way to effectively meet user requirements due to its practical significance (Wan et al., 2020; Zhang et al., 2020; Han et al., 2016).
@@ 4.2 Non-uniform Quantization-based Incremental Upscaling
In contrast to clustering-based methods, state-of-the-art uniform quantization methods often involve complicated mechanisms such as weight reconstruction (Frantar et al., 2023) and per-channel scaling (Lin et al., 2023) that are not readily compatible with incremental upscaling. For example, Table 2 presents the results of applying incremental upscaling to two state-of-the-art uniform methods (AWQ, GPTQ), using a 3-bit model as the seed model. The upscaled models consistently underperform the independently quantized models, showing unacceptable degradation at low bit-widths such as 4. We provide an in-depth discussion on the challenges of applying incremental upscaling on existing uniform quantization methods, along with additional experimental results, in Appendix E.
@@ 5.3 GPU Kernel Optimization
Efficient Bit-transpose. In order to dequantize weights, we have to preform a bit-transpose step (step 2 in Figure 6) as the weights are represented as bitplanes. This step requires a large number of bitwise operations, becoming the main overhead. Even an optimized bit-transpose algorithm requires 38 bitwise operations just to process an 8$\times$8 bit matrix (Warren, 2012).
@@ 6.1 Any-Precision Quantization Results
Methodology. We evaluate 4 to 8-bit models obtained through incremental upscaling, using a 3-bit SqueezeLLM model as the seed model. The results are compared with the 4 to 8-bit SqueezeLLM models. We benchmark our method on LLaMA-2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and three OPT models (6.7B, 2.7B, 1.3B) (Zhang et al., 2022). We evaluate the models with two metrics: perplexity on three datasets (WikiText2 (Merity et al., 2016), PTB (Marcus et al., 1994), C4 (Raffel et al., 2023)) and zero-shot accuracy on five tasks (ARC-easy/challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Tata & Patel, 2003), WinoGrande (Sakaguchi et al., 2021)). Experimental details are in Appendix A.
@@
Table 4: Average zero-shot accuracy on five tasks — ARC-easy, ARC-challenge, HellaSwag, PIQA, and WinoGrande — for vanilla SqueezeLLM (SqLLM) and SqueezeLLM integrated with incremental upscaling (SqLLM+IU) using a 3-bit seed model. We also report the difference between the two methods.
3-bit 4-bit 5-bit 6-bit 7-bit 8-bit FP16
Llama-2-7B SqLLM 66.2 68.3 68.6 68.8 68.9 68.9 69.0
SqLLM+IU - 68.2 68.8 68.9 68.9 69.0
$\Delta$blacktriangledown0.1 $\blacktriangle$0.2 $\blacktriangle$0.1 - $\blacktriangle$0.1
Mistral-7B SqLLM 71.5 73.2 73.8 74.0 74.1 74.2 74.1
SqLLM+IU - 73.3 73.8 74.0 74.2 74.1
[…]
OPT-6.7B SqLLM 58.6 59.9 60.6 60.7 60.7 60.8 60.8
SqLLM+IU - 60.3 60.6 60.7 60.8 60.7
[…]
OPT-2.7B SqLLM 54.3 55.8 56.3 56.2 56.3 56.3 56.4
SqLLM+IU - 55.7 56.1 56.2 56.4 56.3
[…]
OPT-1.3B SqLLM 50.7 52.7 52.9 53.4 53.4 53.3 53.3
SqLLM+IU - 52.6 53.1 53.3 53.4 53.3
[…]
Results. Table 3 shows the perplexity results. The upscaled models (SqLLM+IU) nearly match the independently quantized models (SqLLM) across various models, datasets, and bit-widths. Except in two cases (OPT-6.7B, OPT-1.3B 4-bit on PTB), the increase in perplexity is negligible (< 0.1).
A similar trend is observed in the zero-shot task results. Table 4 presents the average accuracy across five zero-shot tasks. The upscaled models achieve the same level of accuracy as the independently quantized models. The accuracy drop from upscaling is within 0.2%, with upscaled models even outperforming in some cases. Full results for each task are provided in Appendix B.
These results demonstrate that with a single $n$-bit parent model generated through incremental upscaling, we can utilize the full range of 3 to $n$-bit models, all achieving state-of-the-art quality at their respective bit-widths.
Table 5: Runtime (sec) of the proposed any-precision quantization scheme, composed of two stages: seed model generation (Seed Gen) and incremental upscaling (IU).
Llama-2-7B Mistral-7B OPT-6.7B OPT-2.7B OPT-1.3B
Seed Gen 36.2 37.2 37.0 14.0 6.4
IU 15.6 18.2 12.8 6.2 3.0
Total 51.8 55.4 49.8 20.2 9.4
Runtime of Any-Precision Quantization. Our any-precision quantization scheme is efficient as it does not require training and is highly parallelizable. We measure the runtime of the any-precision quantization process, beginning with a 3-bit seed model and progressing up to the final 8-bit parent model, on an Intel i9-13900K CPU with 24 cores. Table 5 shows the results. It takes less than a minute to complete the whole process, even for 7B-scale models.
@@ 6.2 Kernel Microbenchmarks
Methodology. We evaluate the latency of matrix-vector multiplication of our kernel on the three matrix dimensions used in Llama-2-7B, and compare with the existing kernel for non-uniform quantization proposed in SqueezeLLM (Kim et al., 2023b). Note that the SqueezeLLM kernel only supports 3 and 4 bits. We conduct experiments on three GPUs of varying scales: RTX 4090 (desktop), RTX 4070 Laptop (laptop), and Jetson AGX Orin 64 GB (mobile). We use NVIDIA Nsight Compute to measure latency.
Results. Table 6 shows the results. Our kernel consistently achieves low latency for various matrix sizes and platforms, showing a near-linear improvement against the cuBLAS FP16 baseline as the bit-width decreases. Compared to the SqueezeLLM kernel, our kernel performs on par on the RTX 4090 and RTX 4070 Laptop, and goes on to exhibit substantial improvements on Jetson. Note that our kernel achieves the performance despite adopting the bitplane-based weight representation to support any-precision, which entails non-trivial engineering challenges such as addressing the increased amount of bitwise operations for bit-transpose. Additional results with a wider range of matrix dimensions and comparison against uniform quantization kernels can be found in Appendix C.1 and  C.2.
@@
Figure 11: Latency comparison of 4-bit quantized matrix-matrix multiplication using our kernel, separate dequantization followed by cuBLAS FP16 kernel, and cuBLAS FP16 kernel on RTX 4090. Matrix dimensions are ($M$, 4096) and (11008, 4096), with $M$ varying from 1 to 512.
Matrix-Matrix Multiplication Performance. Even for our target setting of running a model on personal devices, where a single query is typically served at a time, efficient matrix-matrix multiplication is necessary for two reasons. First, the batch size might not always be 1; it can be small (e.g., 2, 4, 8). Even in a single query, multiple tokens are generated in parallel when using advanced decoding algorithms like beam search (Sutskever et al., 2014) and parallel sampling. Speculative decoding also requires the parallel processing of multiple tokens. Second, during the prefill phase, the tokens in the input prompt are processed together, with their number potentially reaching into the hundreds or thousands.
The red line in Figure 11 illustrates the latency of our kernel for 4-bit quantized matrix-matrix multiplication with dimensions ($M$, 4096) and (11008, 4096) on an RTX 4090, with $M$ varying from 1 to 512. Our kernel performs robustly with small $M$ values such as 2, 4, and 8, showing a significant latency gap compared to the cuBLAS baseline (green line). This indicates that our kernel is also effective for generation with small batch sizes. Similar results are observed for other bit-widths and platforms, as detailed in Appendix C.3.
When $M$ becomes very large, however, as in the prefill phase, our kernel becomes slower than the cuBLAS FP16 baseline (green line). This slowdown occurs because our kernel does not use tensor cores, making it more susceptible to increases in computation with large $M$. To mitigate potential slowdown of the prefill phase, our engine dequantizes the weights using a separate kernel and then employs the cuBLAS kernel when $M$ exceeds a certain threshold (e.g., 16). The latency gap between the dequantization + cuBLAS (represented by the yellow line) and cuBLAS remains almost constant (about 100 µs), as dequantization time does not scale with $M$. This limited performance degradation of the prefill phase is small enough to be easily amortized when the generation length is sufficiently long.
@@
6.3 End-to-end Throughput
We evaluate the end-to-end inference throughput of our engine by integrating it with TensorRT-LLM (NVIDIA, ). Figure  12 shows the throughput of Llama-2-7B in generating 128 and 1024 tokens on RTX 4090 and RTX 4070 Laptop. TensorRT-LLM currently lacks support for Jetson. The enhanced kernel performance effectively translates into an end-to-end speedup. A slightly lower speedup with a longer sequence length (1024) is due to the increased overhead of attention, which is not accelerated by weight quantization. Results for the other models are in Appendix D.
@@ 7 Conclusion
We make a case for any-precision LLM, which enables memory-efficient and cost-effective deployment of multiple, different-sized LLMs. We propose a lightweight method for any-precision quantization of LLM, along with a specialized software engine to fully leverage its benefits.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea Government (MSIT) (RS-2024-00340008), Institute of Information & Communications Technology Planning & Evaluation (IITP) funded by the Korea Government (MSIT) (2021-0-00105, Development of Model Compression Framework for Scalable On-device AI Computing on Edge Applications), and Institute of Information & Communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program (IITP-2023-RS-2023-00256081), funded by the Korea Government (MSIT).
@@ A.3 Zero-shot Tasks
Appendix B Full Results on Zero-shot Tasks
The data in Table 8 presents the full results for each of the five zero-shot tasks (ARC-easy, ARC-challenge, HellaSwag, PIQA, WinoGrande). The results indicate that the SqLLM models with incremental upscaling (SqLLM+IU) match the independently quantized SqLLM models across a range of tasks, models, and bit-widths.
Appendix C Additional Kernel Microbenchmark Results
C.1 Kernel Latency on Various Matrix Sizes
@@
C.2 Comparison with Kernels for Uniform Quantization
@@
C.3 Matrix-Matrix Multiplication Performance
@@
Appendix D Additional End-to-End Throughput Evaluation Results
@@
Appendix E Incremental Upscaling with Uniform Quantization Methods
@@
In this section, we elaborate on our endeavor to apply incremental upscaling to the two state-of-the-art uniform quantization methods (GPTQ, AWQ), which ultimately results in failure. Figure 14 illustrates the process of incremental upscaling with uniform quantization methods. Upscaling on uniform quantization methods requires that each bin in the quantization grid be divided equally; this is in contrast with non-uniform quantization, where each bin in the quantization grid can be divided into two new bins of arbitrary sizes. Naturally, each bin’s midpoint becomes its representative value. With these premises, Appendix E.1 and  E.2 discuss issues with upscaling GPTQ and AWQ, respectively.
E.1 Incremental Upscaling with GPTQ
@@
E.2 Incremental Upscaling with AWQ